OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8826 — Optic loop Hauts-de-Seine (92)
Incident Report for Web Cloud
Resolved
There was an outage of a few seconds on the NRA optic loop in the 92 area.
Securitisation played its part and the issue has been resolved.
However, an optical signal is damaged, a team is on site to check the
connectors and connections and we have also opened a ticket with the fibre operator.


Update(s):

Date: 2013-06-19 11:27:00 UTC
Operation terminated, the fibre has been repaired. Securitisation is fully operational again.

Date: 2013-06-19 09:33:23 UTC
The operator has found the problem on the fibre optic
and is starting the works.


Date: 2013-06-18 16:53:22 UTC
We have verfied the optic positions on sites DEF92 and
de NAN92, the problem is not on the part that we are responsible for.
We have relaunched the optic operator, and a team is on the way
in order to carry out some reflectometry test on the fibre optics.
Posted Jun 18, 2013 - 13:04 UTC